Use a table to solve. Round to the nearest hundredth.
2^8x = 93

Answer by Alan3354(69443)

Use a table to solve. Round to the nearest hundredth.
2^8x = 93
--------------
8x*log(2) = log(93)
x = log(93)/(log(256)
-----------
What sort of table?